Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Bi-Weekly Remedial Activities Update
Report # 29: March 26 – April 6, 2007

« Back to Atlas Tack Superfund Site Fact Sheet

I. ACTIVITIES PERFORMED DURING THE WEEKS ENDING: 16 March & 23 March

Site Preparation
  • Improved rock apron east of the solidification bins to minimize the tracking of soil over Egypt Lane (large roll-off trucks now using Egypt Lane, Washington St. and Adams St. to access site)
Haul Road Construction (North of Barrier)
  • No road improvements made during this reporting period
  • Mobilized (40) Dura Mats to gain access to re-excavation locations South of the barrier
Excavation
  • Resumed re-excavation activities south of the barrier
  • Began to excavate north of the barrier west of the creek
  • Diverted tributary flowing westward into Boy’s Creek and began to excavate this tributary north of the barrier east of the creek
  • Began to excavate north of the barrier along the toe of the barrier, excavation will continue until hurricane barrier boulders are encountered
Soils Management
  • Continued to consolidate, stockpile, cover and secure stockpiles located in the stockpile staging area
Backfill
  • Began to place imported clean fill in excavation areas located in the marsh north of the barrier east of the creek
Transportation & Disposal
  • Began to transport “contained-in” soils to the TREE landfill facility in Rochester, NH on 3/26/07; operation suspended on 3/27/07
  • Transported sediment to Taunton landfill
  • Waiting for approval of disposal for additional stockpiles from Taunton and TREE
Stabilization
  • All stockpiles treated with the Enviroblend® mixture have been treated successfully:

Transportation & Disposal of Stockpiles

  • CD-019; accepted by TREE
  • CD-021; accepted by TREE
  • WF-016; accepted by Taunton
  • EA-031; accepted by TREE
  • EA-036; accepted by TREE
  • EA-023; accepted by TREE
  • EA-025; accepted by Taunton
  • EA-033
  • EA-034
  • WF-004; accepted by Taunton
  • EA-024
  • EA-026
  • LG-015
De-Watering
  • Continued to handle the upstream waters in Boy’s Creek by using the (3) 6-inch dry prime diesel pumps with a 3-inch trash pump for isolated locations
  • Fairhaven wastewater treatment plant operator reported on 4/3/07 that he would not be able to accept any waters from the site this week due to heavy rainfall
Sampling
  • Collected confirmatory sediment samples from the marsh excavation located north and south of the hurricane barrier
  • Received analytical results from areas of initial excavation north of the barrier and first round of re-excavations south of the barrier
  • Collected (3) clean backfill samples from Marsh Island; results to be provided on 4/4/07
  • Collected (2) samples from clean excavations north of the barrier; results to be provided on 4/6/07
  • Collected (2) sandy loam soil samples from Lopes burrow source, results provided on 4/4/07 and comparison to standards to be performed
Survey
  • Continued to collect post excavation survey limits located north of the hurricane barrier
  • Continued to lay out re-excavation limits
Miscellaneous
  • Continued to monitor both Adams St. and Tripp St entrances and Egypt Lane, inspecting the tractor trailers and removing clinging soil, crossing the trucks over the bike path and sweeping the roads
  • Relocated wood debris over the UST
  • Prepared for standing rainwater discharge

II. POJECTED ACTIVITES FOR THE WEEKS ENDING: 30 March & 6 April

Excavation
  • Continue to excavate north of the barrier, both east and west of Boy’s Creek
  • Continue to address re-excavations identified both north and south of the barrier
Backfill
  • Continue to backfill common fill and sandy loam in areas confirmed to have reached clean-up criteria in both northern and southern sections of the marsh
Soils Management
  • Continue to consolidate, stockpile, cover and secure stockpiles located in the stockpile staging area
Transportation & Disposal
  • Continue to transport COMM-94, COMM-97, COMM-97-Lined, and Sub-D soils/sediment
  • Begin T&D of crushed concrete and brick to Crapo landfill
De-Watering
  • Continue to re-direct normal flow of water in Boy’s Creek and dewater excavation activities
  • Begin to remove standing water in the CID

III. EXCAVATION SUMMARY

WEEKS/E 3/30/07 – 4/6/07
TO DATE
ACTIVITIES
INITIAL TOTAL QUANTITY (cubic yards)
INITIAL TOTAL QUANTITY
Commercial Area Excavation
Complete
2,655 CY
C.I.D. Excavation
None
7,083 CY
East Fill Excavation
None
9,046 CY
NW-Fill Area
None
8,612 CY
Lagoon Area
None
7,683 CY
South of the Barrier
±2,000-cy
±6,888-cy
North of the Barrier
±1,270-cy
±5,875-cy
Total Amount of Soils Excavated to Date:
44,076 - CY

IV. T&D BI-WEEKLY SUMMARY in tons

DATE
DESTINATION
WASTE DETERMINATION
STOCKPILE ID
TONS
4 April 2007
Taunton
COMM-97-Lined
MN-004
±600
3 April 2007
Taunton
COMM-97-Lined
MN-002-003
±600
30 March 2007
Taunton
COMM-97-Lined
MS-012
346.47
29 March 2007
Taunton
COMM-97-Lined
MS-007-012
715.64
28 March 2007
Taunton
COMM-97-Lined
MS-007-012
487.75
27 March 2007
Tree
Sub-D
LG-011-012
926.64
26 March 2007
Tree
Sub-D
EA-008, EA-010, WF-010, LG-010
754.67
21 March 2007
Taunton
COMM-97-Lined
WF-04
379.39

V. T&D TOTAL SUMMARY in tons

WK
DESTINATION
WASTE DETERMINATION
ORIGIN
TONS
6 April 2007
Taunton
COMM-97-Lined
S./N. Barrier
±1,900
30 March 2007
TREE
Sub-D
Lagoon, East/West Fill
1,681.31
30 March 2007
Taunton
COMM-97-Lined
S. Barrier
1,549.86
23 March 2007
Taunton
COMM-97-Lined
S. Barrier, E-Fill
915.8
16 March 2007
Taunton
COMM-97-Lined
S. Barrier, NW-Fill
934.73
16 March 2007
Marion
COMM-97
S. Barrier
280.01
9 March 2007
Marion
COMM-97
S. Barrier
878.46
9 March 2007
TREE
SUB-D
Treated soils from East Fill and soils from south of barrier
1,029.47
2 March 2007
TREE
SUB-D
Treated soils from East Fill
158.04
23 Feb 2007
Marion
SUB-D
Treated soils from East/NW Fill
672.2
16 Feb 2007
Marion
COMM-97
NW/EAST FILL
708.07
9 Feb 2007
Marion
COMM-97
LAGOON/EAST FILL
2054.98
2 Feb 2007
Marion
COMM-97
LAGOON
835.57
26 Jan 2007
Marion
COMM-97
LAGOON
1,322.25
19 Jan 2007
Marion
COMM-97
LAGOON
123.36
12 Jan 2007
Marion
COMM-97
LAGOON
1,314.80
5 Jan 2007
Marion
COMM-97
EAST FILL/NW-FILL
1,458.10
5 Jan 2007
Taunton
COMM-97-Lined
EAST FILL/NW-FILL
201.84
29 Dec 2006
Marion
COMM-97
EAST FILL/NW-FILL
972.76
29 Dec 2006
Taunton
COMM-97-Lined
EAST FILL/NW-FILL
1120.3
22 Dec 2006
Marion
COMM-97
NW-FILL
2343.75
22 Dec 2006
TREE
Sub-D Oversize
EAST FILL
168.6
15 Dec 2006
TREE
Sub-D Oversize
EAST FILL
2,006.51
15 Dec 2006
Marion
COMM-97
NW-FILL
1,134
8 Dec 2006
TREE
Sub-D Oversize
EAST FILL
1,805.81
8 Dec 2006
Marion
COMM-97
NW-FILL/CID
1,143.58
1 Dec 2006
TREE
Sub-D Oversize
EAST FILL
4043.28
24 Nov 2006
TREE
Sub-D Oversize
CID/ EAST FILL
1,590.08
17 Nov 2006
Marion
COMM-97
CID
516.55
17 Nov 2006
Bourne
COMM-97 Oversize
EAST FILL
1,964
10 Nov 2006
TREE
SUB-D
CID
1,695.56
10 Nov 2006
Bourne
COMM-97 Oversize
EAST FILL
274.67
03 Nov 2006
Marion
COMM-97
CID/EAST FILL
2,822.27
27 Oct 2006
T&D DID NOT OCCUR THIS WEEK
20 Oct 2006
T&D DID NOT OCCUR THIS WEEK
13 Oct 2006
T&D DID NOT OCCUR THIS WEEK
06 Oct 2006
Marion
COMM-97
CID
1,264.56
29 Sept 2006
Marion
COMM-97
CID
674.45
22 Sept 2006
Crapo
COMM-97
CID
436.85
15 Sept 2006
Crapo
COMM-97
CID
800.51
T&D TOTAL
45,046.43 tons
30,404.12 CY
 

AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 22 March thru 4 April 2007 can be reviewed in the table below.

Weekly Dust Monitoring Summary
Report 29

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Church Street (µg/m3) Espinolla House Front Yard Espinolla House Rear Stairs
3/22/2007 NSC 0 0 NSC NSC NSC
3/23/2007 NSC 2 0 NSC NSC NSC
3/26/2007 NSC 0 5 NSC NSC NSC
3/27/2007 NSC 0 0 NSC NSC NSC
3/28/2007 NSC 0 0 NSC NSC NSC
3/29/2007 NSC 0 0 NSC NSC NSC
3/30/2007 NSC 0 0 NSC NSC NSC
4/2/2007 NSC NSC NSC NSC NSC NSC
4/3/2007 NSC 0 0 NSC NSC NSC
4/4/2007 NSC 0 0 NSC NSC NSC

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*- Readings are for 24hr period

Jump to main content.